Find m(x) in the interval180(degrees)<(or equal)x >(or equal) 270(degrees) that satisfies the equation 2cos x+1=0
Answer by lwsshak3(6498)

Find m(x) in the interval180(degrees)<(or equal)x >(or equal) 270(degrees) that satisfies the equation 2cos x+1=0
I think you meant 180< x <270 for domain of x
..
2cos x+1=0
2cosx=-1
cosx=-1/2
x= 240 deg in quadrant III